OpenStudy (anonymous):

Solve by using the quadratic formula. Approximate the solutions to the nearest thousandth. x2 − 6x − 5 = 0

OpenStudy (anonymous):

x = [ -b ± sqrt(b^2 - 4ac) ] / 2a from your equation: a=1,b=-6 and c=-5

OpenStudy (anonymous):

yes I got that. And I try and solve it but can't get the answer.

OpenStudy (anonymous):

x=-0.74166 or 6.74166

OpenStudy (anonymous):

now how did u get those answers? last step i knew how to do was.. -6+- sqrt(56) /2

OpenStudy (anonymous):

are you helping me still??

OpenStudy (anonymous):

yes at the last step, you'll obtain a solution like -6+sqrt(56)/2 and -6-sqrt(56)/2

OpenStudy (anonymous):

ok then what do I do next?

OpenStudy (anonymous):

okay, sqrt(56)=7.4833

OpenStudy (anonymous):

so now we have (-6+7.7833)/2 =0.74166

OpenStudy (anonymous):

how is that negative though?

OpenStudy (anonymous):

thats not the answer I got when I divided it by 2

OpenStudy (anonymous):

that last solution will be (-6-7.4833)/2=-6.74166

OpenStudy (anonymous):

and how did u get those? doesnt make any sense to me

OpenStudy (anonymous):

don't divide just the sqrt(56) by 2. it's the whole (-6+sqrt56) divided by 2

OpenStudy (anonymous):

so you take -6+7.7833 and divide that by 2. when I do that I dont get the answer you get.

OpenStudy (anonymous):

and for the second one how did u get 7.4833?

OpenStudy (anonymous):

sqrt(56)=704833

OpenStudy (anonymous):

you mean: 7.4833

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I understand now never mind. but when I put the answers in it is saying that -0.74166 and 6.74166 how is that?

OpenStudy (anonymous):

i'm typing to different people so pardon the typo

OpenStudy (anonymous):

i dont understand how the first one is negative and the second one is positive when i figure it out its the other way around

OpenStudy (anonymous):

your answers will be -0.74166 and 6.74166 because you first the quadratic formula is b plus or minus whatever follows it

OpenStudy (anonymous):

do you mean you got -6.74166 and 0.74166?

OpenStudy (anonymous):

okay a bit confusing.. but thanks now can you help me out with this one: Solve by using the quadratic formula. Approximate the solutions to the nearest thousandth. r2 − 9r = 6

OpenStudy (anonymous):

yes thats what I got

OpenStudy (anonymous):

I think I figured out my problem sorry for all the confusing

OpenStudy (anonymous):

well, it may just be because you you added the 6 instead of (-6)

OpenStudy (anonymous):

okay, the next one, you'll re-arrange to have something like this r2-9r-6=0 and use the same formula

OpenStudy (anonymous):

yes I will tell you what I get for answers and will u tell me if its correct

OpenStudy (anonymous):

okay. i got answers for you when you're done solving

OpenStudy (anonymous):

so far I got 9+sqrt(105)/2 and 9-sqrt(105)/2 is that correct?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

square root of 105 = 10.2470 that right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

14.1235 is that the first answer?

OpenStudy (anonymous):

3.8765 the second answer?

OpenStudy (anonymous):

thats the part where i get stumped

OpenStudy (anonymous):

this what you doing wrong, you supposed to add everything at the top before dividing by 2

OpenStudy (anonymous):

9.6235 that the first answer than?

OpenStudy (anonymous):

so we'll have 9+10.2470=19.247 then we divide 19.247 by 2 to get the correct answer

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

what's your answer for second

OpenStudy (anonymous):

thank you are you willing to help me with one more just so i know i am doing these right..

OpenStudy (anonymous):

-.6235

OpenStudy (anonymous):

is that -0.6235?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me with this one more problem

OpenStudy (anonymous):

just so i know what i am doing

OpenStudy (anonymous):

send

OpenStudy (anonymous):

Solve by using the quadratic formula. Approximate the solutions to the nearest thousandth. 2t2 = 2t + 2

OpenStudy (anonymous):

let me try it all and send u the answers

OpenStudy (anonymous):

3.2361 and -1.2361

OpenStudy (anonymous):

oh crap i know what i did wrong

OpenStudy (anonymous):

1.6180 and -0.6180

OpenStudy (anonymous):

try again. remember to add or subtract the numerators before dividing at the last end

OpenStudy (anonymous):

that's correct. Always ADD or SUBTRACT the top firs before dividing

OpenStudy (anonymous):

i got them now. i divided by 2 which suppose to divided by 4

OpenStudy (anonymous):

ok. take your time and don't rush it unless it's a multiple choice and you have to guess. good luck.